OpenStudy (anonymous):

Simplify quantity 18 x minus 6 over 9 x to the fifth power all over quantity 15 x plus 5 over 21 x squared

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{18x-6}{9x^5}}{\frac{15x+5}{21x^2}}\qquad recall\implies {\color{blue}{ \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}}}\\ \quad \\ \cfrac{18x-6}{9x^5}\cdot \cfrac{21x^2}{15x+5}\implies \cfrac{6(3x-1)}{9x^5}\cdot \cfrac{21x^2}{5(3x+1)}\\ \quad \\ \implies \cfrac{3\cdot 2(3x-1)}{3\cdot 3x^2\cdot x^3}\cdot \cfrac{3\cdot 7x^2}{5(3x+1)}\)

OpenStudy (anonymous):

14x^2 / 15x + 1?

OpenStudy (jdoe0001):

well \(\bf \cfrac{3\cdot 2(3x-1)}{3\cdot 3x^2\cdot x^3}\cdot \cfrac{3\cdot 7x^2}{5(3x+1)}\implies \cfrac{2(3x-1)\cdot 7}{x^3\cdot 5(3x+1)}\\ \quad \\ \implies \cfrac{42x-14}{15x^4+5x^3}\)

OpenStudy (anonymous):

Would that make the answer \[14(3x - 1) \over 5x^3 (3x +1)\] ?

OpenStudy (jdoe0001):

without distributing, yes

OpenStudy (anonymous):

Does it have to be distributed?

OpenStudy (jdoe0001):

no, either way will do

OpenStudy (anonymous):

What would be the distributed form?

OpenStudy (jdoe0001):

ohh ahem well, is the one I posted earlier \(\bf \cfrac{3\cdot 2(3x-1)}{3\cdot 3x^2\cdot x^3}\cdot \cfrac{3\cdot 7x^2}{5(3x+1)}\implies \cfrac{2(3x-1)\cdot 7}{x^3\cdot 5(3x+1)}\\ \quad \\ \implies \cfrac{42x-14}{15x^4+5x^3}\)